OpenStudy (anonymous):

THE amount A that principal P will be worth after t years at interest rate r,compounded annually,is given by this formula A=P(1+r)T suppose 4000 principal is invested at 6% interest compounded annually for five years.how much money would the investment yield after 5 years

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Hi. What specifically would you like to know? This problem deals mostly with substituting known values into a known formula and calculating the amount of $ you'd accumulate.

OpenStudy (anonymous):

i forgot how to solve these problem but are we looking for the interest rate

OpenStudy (mathmale):

Quoting the problem statement: " invested at 6% interest compounded annually " Thus, r=0.06.

OpenStudy (mathmale):

You're not looking for the interest rate; that's given. You're looking to find the amount of money that would accumulate over five years if you invested $4,000 up front and if interest were to be compounded annually (1 time per year).

OpenStudy (anonymous):

ohh do i have to find the amount of p

OpenStudy (mathmale):

p represents "principal" and is the amount you invest up front; that's given in the problem statement: "4000 principal."

OpenStudy (mathmale):

summary: P=principal=$4,000 r=interest rate as a decimal fraction=0.06 t=number of years=5

OpenStudy (anonymous):

would the answer be 1200.

OpenStudy (mathmale):

A=P(1+r)T is not correct; it should be\[A=P(1+r)^t\]

OpenStudy (mathmale):

show me your work, please.

OpenStudy (anonymous):

i multiply 4000 x 0.06 x 5

OpenStudy (anonymous):

to get 1200.00

OpenStudy (mathmale):

Please compare the formula I've given you with what you actually did: \[A=P(1+r)^t\]

OpenStudy (mathmale):

Try again, please.

OpenStudy (anonymous):

4000(1+0.06)5

OpenStudy (anonymous):

21200

OpenStudy (mathmale):

Sorry to have to be picky, but you are not treating the number of years, t, which =5, properly. That t is an exponent. Please start over. I need to see y our work and am not much interested in the "answer."

OpenStudy (mathmale):

Please note that (1+0.06)5 is not the same as (1+0.06)^5, which you should be using.

OpenStudy (anonymous):

4000(1+0.06)^5

OpenStudy (anonymous):

5352.

OpenStudy (mathmale):

Better. Combine 1 and 0.06 to obtain 1.06. Please show how you got that $5,352.

OpenStudy (anonymous):

i used my caculator and put in the equation

OpenStudy (mathmale):

You want A=Amount=$5,000(1.06)^5. what is (1.06)^5?

OpenStudy (anonymous):

1.338

OpenStudy (mathmale):

and, if you multiply that by $5,000, what do you get?

OpenStudy (anonymous):

6691.12

OpenStudy (mathmale):

Right. Better. How did you get 5352. earlier?

OpenStudy (anonymous):

4000(1+0.06)^5

OpenStudy (mathmale):

OK. In this case I'm wrong because I was using $5000 instead of $4000 as the principal. Please, read through our conversation and take notes, so that you'll be better prepared next time to type out the problem and evaluate those exponentials properly.

OpenStudy (mathmale):

Again, (1.06)5 is not the same thing as (1.06)^5. The first one involves multiplication, the second exponentiation, and it's the exponentiation that is necessary here.

OpenStudy (anonymous):

so was i right with my answer of 5352.

OpenStudy (mathmale):

Again I'd prefer that you check your work. (1.06)^5=?

OpenStudy (anonymous):

1.338

OpenStudy (mathmale):

Good; now multiply that by $4,000 (including the $ sign).

OpenStudy (anonymous):

53.52

OpenStudy (mathmale):

You got 1.338 and I asked you to multiply that by $4,000. Where did the 53.52 come from?

OpenStudy (anonymous):

5352

OpenStudy (mathmale):

Better. Please, include the units of measurement. The amount you'll have after 5 years under these terms will be A = ??

OpenStudy (anonymous):

$5352

OpenStudy (mathmale):

Very good. Again, please review our conversation. All of this becomes easier with practice. You've shown that you can use your calculator, but need a lot of prompting along the way. Are we done now?

OpenStudy (anonymous):

yea but im confused ive been saying 5352 in the middle of this conversation and you said it was wrong

OpenStudy (mathmale):

that was because I mistakenly used $5000 as the principal, whereas the correct amount is $4000.

OpenStudy (anonymous):

oh thanks for the help

OpenStudy (mathmale):

The more you show your work, the better I can give y ou accurate feedback. For example: Type out these steps: A=$4000(1+0.06)^5 A=$4000(1.06)^5 A=$4000 (1.338) A=$5352

OpenStudy (mathmale):

thank you for the medal, Panda!